OpenStudy (anonymous):

7x+5y=-3 -9x+y=-11 solve linear system.

OpenStudy (anonymous):

There are two ways of solving a system. You can create a scalar or substitute. Which way did you learn it?

OpenStudy (anonymous):

scalar but i dont get it

OpenStudy (anonymous):

scale?*

OpenStudy (anonymous):

So what you want to do is multiply one of the equations so that one of the variables have additive inverses of the same variable in the other equation( my explanation probably won't make sense. The method with a scalar is called elimination (use this website: http://www.purplemath.com/modules/systlin5.htm)

OpenStudy (anonymous):

7x + 5y = -3 -9x + y = -11 (multiply this equation by 5 so the y's are the same) so it 7x + 5y = -3 -45x + 5y = -55 then subtract the two equations to get rid of the y's so it becomes, 52x = -3 -(-55) 52x = 52 x = 1 sub x in any of the two equations to find y, so 7(1) + 5y = -3 5y= -10 y = -2

OpenStudy (anonymous):

The other way is substitution. You just isolate the y of one of the equations. -9x+y=-11 ---> y=9x-11 and then substitute the y in the other equation 7x+5y=-3 ----> 7x+5(9x-11)=-3 and then just solve the equation to find x. To find y, plug the x in the original equation and solve for y

OpenStudy (anonymous):

ohhh okay got it. thank you guys so much

OpenStudy (anonymous):

someone asked the same question a while ago, same class?

OpenStudy (anonymous):

really? idk

OpenStudy (anonymous):

ya lol exact same,so i just copied pasted the answer here